OpenStudy (anonymous):

solve for x: 4root over 2x-1=1 can anyone help me plzz!

OpenStudy (anonymous):

square both sides to give 16(2x-1)=1 divide by 16 to give 2x-1=1/16 plus one to both sides and then divide by 2 to give x=17/32

OpenStudy (anonymous):

thnkuu so much got it!!